# What’s under the hood: Investigating Automatic Metrics on Meeting Summarization

### v1

@@ Abstract
Meeting summarization has become a critical task considering the increase in online interactions. While new techniques are introduced regularly, their evaluation uses metrics not designed to capture meeting-specific errors, undermining effective evaluation. This paper investigates what the frequently used automatic metrics capture and which errors they mask by correlating automatic metric scores with human evaluations across a broad error taxonomy. We commence with a comprehensive literature review on English meeting summarization to define key challenges like speaker dynamics and contextual turn-taking and error types such as missing information and linguistic inaccuracy, concepts previously loosely defined in the field. We examine the relationship between characteristic challenges and errors by using annotated transcripts and summaries from Transformer-based sequence-to-sequence and autoregressive models from the general summary QMSum dataset. Through experimental validation, we find that different model architectures respond variably to challenges in meeting transcripts, resulting in different pronounced links between challenges and errors. Current default-used metrics struggle to capture observable errors, showing weak to mid-correlations, while a third of the correlations show trends of error masking. Only a subset reacts accurately to specific errors, while most correlations show either unresponsiveness or failure to reflect the error’s impact on summary quality.
@@
Frederic Kirstein1,2, Jan Philip Wahle1, Terry Ruas1, Bela Gipp1 1Georg-August-Universität Göttingen, Germany 2kirstein@gipplab.org
@@ 1 Introduction
With the surge in remote collaboration, there’s an amplified need for proficient meeting summarization Mroz et al. (2018); Pratama et al. (2020), beneficial for participants and non-attendees. Despite advancements in addressing current meeting challenges like contextual turn-taking and discourse structure, existing evaluation metrics often inadequately reflect these Rennard et al. (2023); Kumar and Kabiri (2022). The widespread default use of the ROUGE metric Lin (2004) has shown limitations Liu and Liu (2008); Cohan and Goharian (2016); Fabbri et al. (2021). While the research community acknowledges the potential of newer metrics, such as BERTScore Zhang et al. (2020b), their efficacy in meeting summarization remains uncharted. This is a crucial shortcoming, as the high abstraction level, low extraction rate, and the requirement for complex reasoning power of meeting summarization present new challenges to automatic evaluation metrics, which need to be explored Gao and Wan (2022). Our study aims to investigate how commonly used automatic metrics correlate with human annotations, exploring the inherent qualities these metrics encapsulate in their scores. To unify the fragmented understanding of challenges posed by meeting summarization and the typical errors that manifest when these challenges are unmet. Using the QMSum dataset Zhong et al. (2021), we employ experts to annotate challenges in meeting transcripts and errors in summaries. Various models, including domain-standard encoder-decoder architectures and notable autoregressive models, generated these summaries. Our methodological approach enables us to draw correlations between the challenges and errors, the annotated errors, and a set of eight automatic metrics. The insights derived reveal shortcomings in the current automatic metrics employed for meeting summarization evaluation: Structural disorganization errors are often penalized, reflecting human assessments, whereas hallucination errors tend to be rewarded. While adept at penalizing omissions, the ROUGE Lin (2004) metric struggles to discern the nuances in error impact on quality. Strikingly, around a third of metrics-error combinations either overlook or reward errors. For instance, Perplexity leans towards incorrect references, and Lens tends to favour structural disorganization. This underscores a pressing demand for refined evaluation methods in meeting summarization. Our contributions are threefold:
@@
  • We build the first direct correlations between the intrinsic challenges and observable errors they induce for encoder-decoder and autoregressive model architectures, providing a framework for tracking their impact.
@@
We will release the codebases, annotations and guidelines to aid continued research in this domain.
@@ 2 Methodology
We conducted a focused literature review to identify challenges and observable errors when failing these challenges in English meeting summarization, specifically targeting papers published after the seminal 2019 introduction of the transformer-based BART Lewis et al. (2020) model, which left a lasting mark in the field. This further allows us to target the challenges in meeting summarization with our survey as they emerge and pertain to the currently used Transformer-based models (e.g., coreference resolution and spoken language). Besides this temporal filter, we set inclusion criteria to prioritize relevance over historical comprehensiveness and consider works that either introduce novel methodologies or contribute to defining challenges, with a preference for those employing transformer architectures. Exclusion criteria eliminate tangential studies, highly specialized papers not addressing broader challenges, multi-lingual and multi-modal approaches, and duplicates. We considered publications from 2023 irrespective of their citation index, acknowledging their emergent influence. The survey thereby orients on the PRISMA checklist to prevent bias.
Our primary database was Google Scholar , chosen for its comprehensive coverage of both extractive and abstractive methodologies and advanced query capabilities. While ACL Anthology , Web of Science , and DBLP were assessed, Google Scholar yielded a more nuanced spectrum of relevant literature, conforming to our review scope spanning from 2019 to September 2023.
A two-stage search strategy was employed: The initial phase involved broad queries combining "meeting summarization" with adjunctive terms such as challenge, literature review, or survey, combined through a logical operation. Subsequent refinement employed challenge-related keywords and synonyms retrieved from the review literature, ensuring alignment with our research focus. Papers were subsequently ranked based on their Google Scholar metrics, with the top 100 from each query considered. The corpus thus aggregated served as the primary source for our analysis.
From the initial pool of 300 papers, we rigorously applied our criteria to select 18 core articles addressing challenges directly and an additional 40 that discuss challenges and observable errors indirectly through methodological contributions. Each shortlisted paper underwent detailed data extraction to identify relevant challenges, methodologies, foundational models, and metrics, which were then collated in a structured format for cross-validation. This allowed us to construct a comprehensive view of current challenges and observable errors in English meeting summarization.
3 Literature Review
@@ 3.1 Challenges
Spoken language. Handle colloquialisms, domain-specific terminology, and various forms of linguistic noise, such as false starts, repetitions, and filler words Koay et al. (2020); Kumar and Kabiri (2022); Feng et al. (2022); Antony et al. (2023); Zhang et al. (2021), which can appear due to the spoken language nature of meetings. This can affect the accuracy and clarity of the generated summaries.
Speaker dynamics. Accurately distinguish and track different speakers, their utterances, and specific roles (e.g., project manager, applicant), particularly when roles are topic-dependent Gu et al. (2022); Rennard et al. (2023); Shinde et al. (2022); Feng et al. (2022); Antony et al. (2023); Kumar and Kabiri (2022); Khalifa et al. (2021). Failing to do so can introduce biases and result in incomplete or flawed summaries leaving out noteworthy elements.
Coreference. Manage the resolution of references to other speakers and previous actions to ensure coherent and complete summarization Liu et al. (2021). Inaccurate handling can lead to ambiguous or incomplete summaries lacking contextual richness.
Discourse structure. Understand and track the inherent high-level structure and flow of a meeting Li et al. (2023) throughout the different meeting phases. These phases may refer to multiple topics Feng et al. (2022), e.g., during an argumentation, or swiftly shift from one topic to another Shinde et al. (2022); Feng et al. (2022). Failure to consider this hierarchical structure can result in summaries that are either incomplete or lack coherence.
Contextual turn-taking. Capture the evolving local dynamics of a meeting as it progresses through multiple speaker turns. The task involves accurately processing shifts in discourse, complicated by interruptions, repetitions, and redundancies Ma et al. (2022); Gu et al. (2022); Zhang and Zhao (2021); Zhang et al. (2021); Shinde et al. (2022). Inadequate capture can result in misleading or shallow summaries.
@@
Low information density. Identify key points when salient information is sparse and unevenly distributed, especially relevant in decision-making scenarios Rennard et al. (2023); Zhang et al. (2021). This can affect the depth and level of salience in the resulting summary.
Data scarcity. The lack of diverse, high-quality real-world scenario training samples hampers model training Rennard et al. (2023); Kumar and Kabiri (2022); Jacquenet et al. (2019).
Long transcripts. Process long transcripts, which can result from longer meetings, due to the quadratic computational cost of transformer-based models, leading to efficiency issues with increasing dialogue length Rennard et al. (2023); Kumar and Kabiri (2022); Shinde et al. (2022); Feng et al. (2022); Jacquenet et al. (2019); Zhang et al. (2021). Only processing a sub-part of a meeting might lead to the exclusion of salient information and therefore result in an incomplete or wrong summary.
Heterogeneous meeting formats. Different types of meetings necessitate distinct summarization approaches Rennard et al. (2023). Failure to adapt can result in summaries that miss key points or include irrelevant information
@@ 3.2 Error Types
Errors in summaries arise when challenges aren’t correctly addressed. We’ve collated error types from the literature into six principal categories, which were observed in the summaries generated during our experiments as detailed in Section 5.
Missing information. This error involves missing information from the meeting, such as significant decisions or actions Zou et al. (2023); Chen and Yang (2020a). We consider total omission, where relevant topics or points discussed are entirely absent from the summary, and insufficient detail, where the summary mentions a salient topic but does not capture its depth or the detailed discussion from the original meeting.
Redundancy. The summary contains repeated or redundant information, impacting brevity and clarity Chen and Yang (2020a). Such repetitions can manifest in different ways: through reiterated key points, overuse of individual words, or the duplication of entire phrases.
Wrong references. The model misattributes statements, opinions, or actions to incorrect meeting participants or omits their mention altogether Chen and Yang (2020a).
Incorrect reasoning. The model draws conclusions that are not supported by the discussions in the meeting Chen and Yang (2020a).
Hallucination. The model produces inconsistencies, such as incorrect dates, names, or locations, not aligned with the meeting content Wang et al. (2022a); Ji et al. (2023). This encompasses intrinsic hallucinations, introducing events not present in the original meeting or contradicting the input, and extrinsic hallucinations, misrepresenting actual events.
Incoherence. The model generates summaries with disjointed logic or flow. This manifests as intra-sentence disconnections and inter-sentence inconsistencies, such as flawed progression and incorrect contrastive expressions Wang et al. (2023).
During our human annotation process, as detailed in Section 4.4, we identified two additional error categories. Notably, our review of the pertinent literature revealed that these categories have not been explicitly defined in existing studies in a way that matches our findings.
Linguistic inaccuracy. The model uses inappropriate, incorrect, or ambiguous language or fails to capture unique linguistic styles. This error type spans issues from adopting unsuitable words from the source to grammar mistakes, as well as employing contextually unclear or ambiguous terms.
Structural disorganization. The model may produce summaries that misrepresent the original order or logic of the meeting’s discourse, misplacing topics or events. This error covers only when the order of stated events is wrong. The inclusion of non-existing events or exclusion of events does not count towards this error.
@@ 4.1 Models
Our primary model is the Longformer Encoder Decoder (LED) Beltagy et al. (2020) given its recognized summarization task performance. LED employs a unique local+global sliding window attention in its encoder and full self-attention in its decoder, accommodating lengthy documents efficiently. For a comprehensive assessment of challenges and errors in the context of the common encoder-decoder architecture, we incorporated DialogLED Zhong et al. (2022) and PEGASUS-X Phang et al. (2022). DialogLED extends LED with dialogue-centric pre-training. In contrast, PEGASUS-X enhances the original PEGASUS Zhang et al. (2020a) for extended inputs using staggered block-local attention. These models have been finetuned on QMSum.
We also ventured into large autoregressive language models for meeting summarization, leveraging GPT3.5 turbo via ChatGPT and Zephyr-7B-$\alpha$, a refined version of Mistral-7B-v0.1111https://huggingface.co/mistralai/Mistral-7B-v0.1. Mistral’s adaptations employ sliding window attention and have gained recent attention for outperforming Llama2 Touvron et al. (2023). These models, exemplifying the trends in language models, excel in zero-shot summarization Bhaskar et al. (2023). To tackle context size limitations, we employed a chunk-based approach Bhaskar et al. (2023), using the prompt "Create a TL;DR of the following meeting chunk", inspired by a Microsoft guideline  [^2] https://www.microsoft.com/en-us/microsoft-365-life-hacks/organization/using-chatgpt-creating-meeting-agendas-minutes-notes.
@@ 4.2 Datasets
We employ QMSum Zhong et al. (2021), a specialized, widely established, and currently most comprehensive dataset for query-based multi-domain meeting summarization. It includes a diverse range of meeting transcripts from academic meetings (ICSI), product meetings (AMI), and committee meetings (Welsh Parliament, Canadian Parliament: WPCP). Detailed statistics are available in Table 1.
ICSI Janin et al. (2003) offers transcripts from informal research meetings, showcasing linguistic challenges such as missing words, sentence restarts, and fillers. AMI Mccowan et al. (2005) provides structured meetings with predefined participant roles and agendas yet still captures natural dialogue dynamics like interruptions and non-verbal cues. WPCP presents formal, agenda-driven discussions from UK and Canada Parliament committee meetings, allowing for a more structured summarization task.
SL SD Cor DS CTT IC DD LID overall
challenges 0.84 0.77 0.68 0.79 0.73 0.90 0.87 0.78 0.79
@@
Encoder-Decorder 0.74 0.875 0.83 0.87 0.79 0.61 0.86 0.71 0.78
impact (LED) 0.82 0.92 0.85 0.73 0.80 0.89 0.87 0.79 0.83
Autoregressive 0.76 0.87 0.85 0.79 0.77 0.90 0.90 0.71 0.82
Table 3: Krippendorff’s alpha for inter-annotator agreement on errors for encoder-decoder (i.e., LED, DialogLED) and autoregressive models (i.e., GPT-3.5, Zephyr-7B-$\alpha$). For the LED model, the agreement on error impact is also reported. Errors (abbreviations) in the top row are ordered as in Section 3.2.

### v2

@@ Abstract
Meeting summarization has become a critical task considering the increase in online interactions. Despite new techniques being proposed regularly, the evaluation of meeting summarization techniques relies on metrics not tailored to capture meeting-specific errors, leading to ineffective assessment. This paper explores what established automatic metrics capture and the errors they mask by correlating metric scores with human evaluations across a comprehensive error taxonomy. We start by reviewing the literature on English meeting summarization to identify key challenges, such as speaker dynamics and contextual turn-taking, and error types, including missing information and linguistic inaccuracy, concepts previously loosely defined in the field. We then examine the relationship between these challenges and errors using human-annotated transcripts and summaries from encoder-decoder-based and decoder-only Transformer models on the QMSum dataset. Experiments reveal that different model architectures respond variably to the challenges, resulting in distinct links between challenges and errors. Current established metrics struggle to capture the observable errors, showing weak to moderate correlations, with a third of the correlations indicating error masking. Only a subset of metrics accurately reacts to specific errors, while most correlations show either unresponsiveness or failure to reflect the error’s impact on summary quality.
@@
Frederic Kirstein1,2, Jan Philip Wahle1, Terry Ruas1, Bela Gipp1 1University of Göttingen, Germany 2kirstein@gipplab.org
@@ 1 Introduction
The rise in remote collaboration has increased the need for effective meeting summarization Mroz et al. (2018); Pratama et al. (2020), beneficial for participants and non-attendees. However, established evaluation metrics do not fully capture the challenges of meeting transcripts such as contextual turn-taking and discourse structure Rennard et al. (2023); Kumar and Kabiri (2022). The commonly used ROUGE metric Lin (2004) has limitations in reflecting summary quality Liu and Liu (2008); Cohan and Goharian (2016); Fabbri et al. (2021), and newer metrics like BERTScore Zhang et al. (2020b) have not been thoroughly tested for meeting summarization Kirstein et al. (2024). Yet, this assessment would be crucial because meeting summarization comes with unique challenges, such as high abstraction, low extraction rate, and complex reasoning Gao and Wan (2022).
This study examines how automatic metrics relate to human annotations for meeting summarization and what they actually measure in their scores. We aim to create a unified understanding of the challenges in meeting summarization and the errors that occur when these challenges are unmet. Using the QMSum dataset Zhong et al. (2021), we have experts annotate the challenges in the transcripts and errors in automatically generated summaries using various models, including domain-standard encoder-decoder architectures and notable decoder-only models. This setup allows us to find connections between challenges, errors, and eight automatic metrics.
The results show problems with current automatic metrics for evaluating meeting summarization. Structural disorganization errors are often penalized, matching human judgments, but hallucination errors are sometimes rewarded. ROUGE Lin (2004) struggles to distinguish the impact of different errors on quality, even though it is good at penalizing omissions. Surprisingly, about a third of the metrics-error combinations either ignore or reward errors. For example, Perplexity favors incorrect references, and Lens favors structural disorganization. These observations highlight the need for better evaluation methods in meeting summarization. Our contributions are threefold:
@@
  • We build the first direct correlations between the intrinsic challenges and observable errors they induce for encoder-decoder and decoder-only model architectures, providing a framework for tracking their impact.
@@
The codebase, annotations, and guidelines are available on GitHub:
@@ 2 Methodology
We conducted a focused literature review to identify challenges and observable errors when failing these in English abstractive meeting summarization. We target papers published between the introduction of the BART model Lewis et al. (2020) in 2019, which serves as the typical backbone model for the field, and 2024. This scope allows us to focus on challenges pertaining to currently used Transformer-based models. We set additional inclusion criteria to prioritize relevance. We consider works that introduce novel methodologies or contribute to defining challenges, with a preference for transformer architectures, and exclude only loosely related studies, multi-lingual, multi-modal approaches, non-abstractive, and non-Transformer-based approaches. We considered publications from 2024, irrespective of their citation index, acknowledging their emergent influence. The survey follows the PRISMA checklist Page et al. (2021) to prevent bias.
Google Scholar is chosen as the primary database for its comprehensive coverage and advanced query capabilities, yielding more works conforming to our review scope than Web of Science and DBLP.
Our search comprises two stages: The initial phase involves broader queries combining "meeting summarization" with adjunctive terms such as challenge, literature review, or survey, while the subsequent refinement uses challenge-related keywords and synonyms. Papers are ranked based on Google Scholar metrics, considering the top 100 from each query, and serve as primary source for our analysis.
We apply our criteria to the initial pool of 300 papers to select 18 core articles directly addressing challenges and 40 additional papers discussing challenges and errors implicitly through methodological contributions. Each shortlisted paper undergoes detailed data extraction to identify relevant challenges, methodologies, foundational models, and metrics collated in a structured format for cross-validation. This approach allows us to construct a comprehensive view of current challenges and observable errors in English meeting summarization.
3 Definitions
@@ 3.1 Challenges
Spoken language. Handle colloquialisms, domain-specific terminology, and various forms of linguistic noise, such as false starts, repetitions, and filler words Koay et al. (2020); Kumar and Kabiri (2022); Antony et al. (2023), which can appear due to the spoken language nature of meetings. This challenge can affect the accuracy and clarity of the generated summaries.
Speaker dynamics. Accurately distinguish and track different speakers, their utterances, and specific roles (e.g., project manager, applicant), particularly when roles are topic-dependent Khalifa et al. (2021); Gu et al. (2022); Rennard et al. (2023). Failing to do so can introduce biases and result in incomplete or flawed summaries, leaving out noteworthy elements.
Coreference. Manage the resolution of references to other speakers and previous actions to ensure coherent and complete summarization Liu et al. (2021). Inaccurate handling can lead to ambiguous or incomplete summaries lacking context.
Discourse structure. Understand and track the inherent high-level structure and flow of a meeting Li et al. (2023) throughout the different meeting phases. These phases may refer to multiple topics Feng et al. (2022), e.g., during an argumentation, or swiftly shift from one topic to another Shinde et al. (2022). Failure to consider this hierarchical structure can result in summaries that are either incomplete or lack coherence.
Contextual turn-taking. Capture the evolving local dynamics of a meeting as it progresses through multiple speaker turns. The task involves accurately processing shifts in discourse complicated by interruptions, repetitions, and redundancies Ma et al. (2022); Zhang and Zhao (2021); Shinde et al. (2022). Inadequate capture can result in misleading or shallow summaries.
@@
Low information density. Identify key points when salient information is sparse and unevenly distributed, especially relevant in decision-making scenarios Rennard et al. (2023); Zhang et al. (2021). This challenge can affect the resulting summary’s depth and level of salience.
Data scarcity. The lack of diverse, high-quality, real-world scenario training samples hampers model training Rennard et al. (2023); Kumar and Kabiri (2022); Jacquenet et al. (2019).
Long transcripts. Process long transcripts, which can result from longer meetings due to the quadratic computational cost of transformer-based models, leading to efficiency issues with increasing dialogue length Kumar and Kabiri (2022); Feng et al. (2022); Zhang et al. (2021). Only processing a sub-part of a meeting might exclude salient information and, therefore, result in an incomplete or incorrect summary.
Heterogeneous meeting formats. Different types of meetings necessitate distinct summarization approaches Rennard et al. (2023). Failure to adapt can result in summaries that miss key points or include irrelevant information.
@@ 3.2 Error Types
Errors in summaries arise when challenges are not correctly addressed. We have collated error types from the literature into six principal categories, observed in the summaries generated during our experiments as detailed in Section 5.
Missing information (MI). This error involves missing information from the meeting, such as significant decisions or actions Zou et al. (2023); Chen and Yang (2020). We consider total omission, where relevant topics or points discussed are entirely absent from the summary, and insufficient detail, where the summary mentions a salient topic but does not capture its depth or the detailed discussion from the original meeting.
Redundancy (Red). The summary contains repeated or redundant information, impacting brevity and clarity Chen and Yang (2020). Such repetitions can manifest in different ways: reiterated key points, overuse of individual words, or duplicating entire phrases.
Wrong references (WR). The model misattributes statements, opinions, or actions to incorrect meeting participants or omits their mention altogether Chen and Yang (2020).
Incorrect reasoning (IR). The model draws conclusions that are not supported by the discussions in the meeting Chen and Yang (2020).
Hallucination (Hal). The model produces inconsistencies, such as incorrect dates, names, or locations, not aligned with the meeting content Wang et al. (2022); Ji et al. (2023). This encompasses extrinsic hallucinations, introducing events not present in the original meeting or contradicting the input, and intrinsic hallucinations, misrepresenting actual events.
Incoherence (Inc). The model generates summaries with disjointed logic or flow. This manifests as intra-sentence disconnections and inter-sentence inconsistencies, such as flawed progression and incorrect, contrastive expressions Wang et al. (2023).
We have identified two additional error categories during our human annotation process, as detailed in Section 4.3. Notably, our review of the pertinent literature reveals that these categories have not been explicitly defined in existing studies in a way that matches our findings.
Linguistic inaccuracy (LI). The model uses inappropriate, incorrect, or ambiguous language or fails to capture unique linguistic styles. This error type spans issues from adopting unsuitable words from the source to grammar mistakes and employing contextually unclear or ambiguous terms.
Structural disorganization (SD). The model may produce summaries that misrepresent the original order or logic of the meeting’s discourse, misplacing topics or events. This error covers only when the order of stated events is wrong. Including non-existing events or excluding events does not count towards this error.
@@ 4.1 Models
We use the Longformer Encoder Decoder (LED) Beltagy et al. (2020) as our primary model for its strong summarization performance. LED employs local+global sliding window attention in the encoder and full self-attention in the decoder, efficiently handling lengthy documents. To comprehensively assess challenges and errors in encoder-decoder architectures, we further consider DialogLED Zhong et al. (2022) and PEGASUS-X Phang et al. (2022). DialogLED extends LED with dialogue-centric pre-training, while PEGASUS-X enhances PEGASUS Zhang et al. (2020a) for extended inputs using staggered block-local attention. All models are finetuned on the QMSum training subset (general summaries) for three epochs.
We also explore large language models, using GPT3.5 turbo via ChatGPT and Zephyr-7B-$\alpha$, a refined version of Mistral-7B-v0.1  [^1] https://huggingface.co/mistralai/Mistral-7B-v0.1. Mistral’s adaptations employ sliding window attention, outperforming Llama2 Touvron et al. (2023). To handle context size limitations, we use a chunk-based approach Bhaskar et al. (2023) with the prompt "Create a TL;DR of the following meeting chunk," inspired by a Microsoft guideline  [^2] https://www.microsoft.com/en-us/microsoft-365-life-hacks/organization/using-chatgpt-creating-meeting-agendas-minutes-notes. LLMs are used with a zero-shot setup.
For the experiments, the models are grouped into encoder-decoder models containing LED, DialogLED, and Pegasus-X, and the decoder-only models, i.e., GPT-3.5 turbo and Zephyr-7B-$\alpha$. This grouping is motivated by similar architecture and error distribution and frequency patterns as shown in Table 8.
@@ 4.2 Datasets
We use as input samples from QMSum Zhong et al. (2021), an established dataset for query-based multi-domain meeting summarization. It includes transcripts from academic (ICSI), product (AMI), and committee meetings (Welsh and Canadian Parliament: WPCP). ICSI Janin et al. (2003) offers informal research meetings with linguistic challenges. AMI Mccowan et al. (2005) provides staged meetings with natural dialogue dynamics. WPCP presents formal, agenda-driven discussions from UK and Canada Parliament committee meetings. Detailed statistics are in Table 1. While there are other datasets out there in the context of meeting summarization, such as the frequently used MeetingBank Hu et al. (2023) and ELITR Nedoluzhko et al. (2022) datasets, we do not include these. MeetingBank is conceptually close to the WPCP meetings, therefore not enhancing the diversity of our data selection and potentially introducing an imbalance in meeting type. ELITR aims to produce meeting minutes with bullet points containing key insights and actions discussed. Therefore, it does not fit the abstractive summaries we aim to investigate.
Spoken Language Speaker Dynamic Co- reference Discourse Structure Contextual Turn-Taking Implicit Context Data Scarcity Low Inf. Density Overall
detection 0.84 0.77 0.68 0.79 0.73 0.90 0.87 0.78 0.79
@@
detection Encoder-Decorder 0.74 0.875 0.83 0.87 0.79 0.61 0.86 0.71 0.78
Decoder-only 0.76 0.87 0.85 0.79 0.77 0.90 0.90 0.71 0.82
error impact (LED) 0.82 0.92 0.85 0.73 0.80 0.89 0.87 0.79 0.83
Table 3: Krippendorff’s alpha for inter-annotator agreement on errors for encoder-decoder (i.e., LED, DialogLED) and decoder-only models (i.e., GPT-3.5, Zephyr-7B-$\alpha$). For the LED model, the agreement on error impact is also reported. Errors (abbreviations) in the top row are ordered as in Section 3.2.